The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below: Standard Hours Standard Rate Standard Cost Motor tune-up 2.50 $29.00 $72.50 The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 54 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups: Labor rate variance $ 120 F Labor spending variance $ 170 U Required: 1. Determine the number of actual labor-hours spent on tune-ups during the week. 2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.
Answer #1
Labor Efficiency variance = Labor rate variance+Labor spending variance
Labor Efficiency variance = 120+170 = $ 290 U
1
Labor Efficiency variance = SR*(AH-SH)
290 = 29.00*(AH-54*2.50)
10 = AH-135
AH = 145
Actual labor-hours = 145 hours
2
Labor rate variance = AH*(SR-AR)
120 = 145*(29.00-AR)
0.83 = 29.00-AR
AR = 28.17
Actual hourly rate of pay = 28.17 per hour